5.1 Assume that a company is negotiating to do business with a new supplier who has offered credit terms of 3/15, net 30. The financial manager is planning to delay payment for an additional 10 days i.e. to only settle the account after 40 days. The current bank overdraft rate for the firm is 25% per annum. Calculate the effective cost of finance provided by this supplier and comment on the financial manager’s plans. (5) 5.2 An aspect of working capital policy which requires managerial attention is the manner in which the items are financed.

Required:

Discuss aggressive, moderate and conservative policies in this regard
Answer #1

Cost of not Discount% taking a cash=- -X- discount 100%-Disc.% 360 final due date- discount period

= 3% / (100% - 3%) x 360 / (30 - 15) = 74.23%

The effective cost of finance provided by this supplier = 74.23%

Manager's plan: He will delay the payment by further 10 days. In that case the final due date becomes = 30 + 10 = 40 days.

Cost of not taking discount then becomes = = 3% / (100% - 3%) x 360 / (40 - 15) = 44.54% > 25% = the bank's overdraft rate. Hence, the manager's plan is not optimal. Manager should borrow and pay the supplier on 15th day and enjoy the discount.

Aggressive policy: Short term funds are used to finance the long term assets. Example will be: Usage of short term loans / working capital loans / overdraft to finance the long term fixed assets like plant, machinery, equipment etc. This results into asset liability maturity mismatch.

Moderate: Short term funds are used to finance short term assets and long terms funds are used to finance the long term assets. This is most widely followed working capital management policy. This leads to usage of working capital loans / revolving facility / overdraft towards funding / financing receivables and or inventory.

Conservative: Usage of long term funds towards short term assets. Company borrows on long term basis to finance the working capital needs. This is sub optimal.